Assuming that the probability of a pea having a green pod is 0.15. When 12 peas are generated, use Binomial probability formula to find the probability of getting at least 2 green peas.
Answer by SUNIL1987MAURYA(7)

probability of a pea having a green pod is= 0.15
probability of a pea having not a green pod is=1-0.15=0.85
the probability of not getting at least 2 green peas=(12)C(0)(0.15)^0(0.85)^12+(12)C(1)(0.15)^1(0.85)^11
=(0.85)^11(0.85+12*0.15)
=0.443346
therefore the probability of getting at least 2 green peas=1-0.443346=0.556654